OpenStudy (anonymous):

which equation is solved for the variable c? a(c+b)=d c= d-ab/a c=d+ab/a c=ab-d/a c=d-ab/b

Vocaloid (vocaloid):

start with a(c+b)=d now divide both sides by a

OpenStudy (anonymous):

c=-b+d/a?

Vocaloid (vocaloid):

excellent!

OpenStudy (anonymous):

how do i find which answer choice that would be?

OpenStudy (jhannybean):

a(c+b)=d ac +ab = d ac = d - ab Now try solving for c.

OpenStudy (anonymous):

c=d/a-b?

OpenStudy (anonymous):

i think its A

OpenStudy (jhannybean):

Correct.

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

Np